# Security Weaknesses of Copilot Generated Code in GitHub

### v1

@@ 5.2. Implications
Code Snippets with Security Weaknesses: In practical production, practitioners often use Copilot to generate code in six languages: Python, JavaScript, Java, C++, Go, and C#. These languages all inevitably produce security weaknesses. We conjecture that practitioners using Copilot will likely encounter security weaknesses, regardless of the programming language used, and security checks are mandatory. When using Copilot, practitioners should conduct their own assessment of the generated code with the support of security analysis tools. They should exercise extreme caution when attempting to rely entirely on Copilot’s behavior, especially for the most commonly used languages with Copilot: Python and JavaScript.
Types of Security Weaknesses in Copilot Generated Code: Practitioners using Copilot may encounter a variety of security weaknesses. Our results show that these weaknesses are related to over 40 CWEs. This finding indicates that there are diverse security scenarios in production, and practitioners must have the corresponding security awareness and skills and adopt multiple security prevention measures to address security risks so that they do not simply accept vulnerable code suggestions. At the same time, our study reveals the frequency of related CWEs. When using Copilot to generate code, practitioners should pay particular attention to specific weaknesses, such as CWE-78: OS Command Injection, CWE-330: Use of Insufficiently Random Values, CWE-502: Deserialization of Untrusted Data, CWE-703: Improper Check or Handling of Exceptional Conditions, and CWE-400: Uncontrolled Resource Consumption. Our findings can assist practitioners in proactively preventing and addressing security issues in a targeted manner.
The CWEs in Copilot Generated Code from the Top-25 CWE List: Common security weaknesses in software development are also prevalent in code generated by Copilot. As a good practice, developers can use the CWE Top-25 list as a guide to understand which security weaknesses are most common and dangerous in the generated code and take appropriate measures to improve the code security. Additionally, the CWE Top-25 provides a standardized approach for security assessment, and developers can also use it to conduct security audits of the code generated by Copilot. Developers should also follow best practices and use code analysis tools (static, dynamic, or hybrid) to check the suggested code by Copilot (or any code generation tools) before integrating any code suggestions. Such tools can safeguard the code and help in discovering weaknesses early.
@@ 6. Threats to Validity
Construct Validity is the degree to which a measurement can explain the theoretical structure and characteristics of the measurement, reflecting the extent to which the studied operational measures truly represent the researcher’s ideas and the content investigated based on the research questions. This study has three threats to the construct validity: (1) Using the keyword-based search – We used a keyword-based search to collect relevant code snippets from GitHub. The results obtained through the keyword-based search may not cover all code snippets generated by Copilot on GitHub. We tried to mitigate this threat by constantly and iteratively refining the keywords and using synonyms. (2) Manual data filtering – We manually screened the results obtained from the keyword-based search by analyzing the comments, tags, and other metadata of the code snippets to determine whether they were generated by Copilot. Since this process was manually done, it may have been influenced by personal bias. At the same time, we assumed that all code files contained in projects declared to be written by Copilot in markdown files were generated by Copilot, and we included them in the research data. This has an impact on the construct validity, as we could not have excluded the possibility of human-written code files in these projects. (3) Manual association of CWEs – We manually associated some warning messages prompted by static analysis tools with CWEs. Some warning messages could be associated with multiple CWEs, while we only focused on assigning one most suitable CWE for each warning message. Personal bias may occur with this step, impacting the final association, and we mitigated the bias by two authors conducting the association with the assessment by a security expert.
@@
External Validity refers to the extent to which research results can be generalized and the degree to which people outside the investigated cases are interested in the research results. It indicates whether the research results are representative and can be validated in similar contexts. Our dataset consists of Copilot generated code snippets collected from open-source projects on GitHub. During the filtering process, we excluded code that utilized Copilot to solve algorithmic problems, aiming to ensure that the collected data genuinely reflected real-world production environments. Due to the reason that the data from GitHub is not diversified enough, there are many code snippets from Game projects. The peculiarity of the data source may make the dataset incomplete, thereby threatening the external validity of the results. However, GitHub is one of the largest code hosting platforms in the world, with hundreds of millions of public code repositories, and is popular among developers and the technology community. The code snippets obtained from GitHub are diverse, which mitigates this threat. Furthermore, we acknowledge the need to collect more diverse code snippets from different platforms to increase the generalizability of the results. We will consider adopting more diversified ways or platforms to collect code. Additionally, due to the limitations of static analysis tools themselves, these tools could not scan all CWEs, and there is a degree of false-positives in the scanning results (as the case with static analysis tools (Sui et al., 2020; Kang et al., 2022)). Although we attempted to use two static analysis tools to increase the coverage of weaknesses, these tools may have limited abilities in analyzing some CWEs with specific error rates, which may impact the completeness and correctness of the results.
@@
Reliability refers to the extent to which a specific research method can produce consistent results. We used multiple automated static analysis tools to analyze the Copilot generated code snippets to improve security weaknesses detection. Developers have widely used these automated tools. The querying mechanism of these tools ensures that the scan results remain consistent when used multiple times. In addition, we performed two rounds of scanning with two tools for security checks on each code snippet, intending to complement the results of one tool with the other one. By implementing these measures, we believe that our research results are reliable and these threats to reliability are mitigated.
@@ 7. Conclusions
Automatic code generation and recommendation has been an active research area due to the advancement of AI and specifically LLMs. AI code generation tools such as Copilot can greatly improve the development efficiency of programmers, but they can also introduce vulnerabilities and security risks. In this paper, we present the results of an empirical study to analyze security weaknesses in Copilot generated code found in public GitHub projects. We identified 435 code snippets generated by Copilot from GitHub projects and analyzed those snippets for security weaknesses using static analysis tools. This study aims to help developers understand the security risks of weaknesses introduced in the code generated by Copilot (and potentially similar code generation tools). Our results show: (1) 35.8% of the 435 Copilot generated code snippets contain security weaknesses, spreading across six programming languages. (2) The detected security weaknesses are diverse in nature and are associated with 42 different CWEs. The CWEs that occurred most frequently are CWE-78: OS Command Injection, CWE-330: Use of Insufficiently Random Values, and CWE-703: Improper Check or Handling of Exceptional Conditions (3) Among these CWEs, 11 appear in the MITRE CWE Top-25 list, demonstrating their commonality, current and severity. These are: CWE-78: OS Command Injection, CWE-502: Deserialization of Untrusted Data, CWE-400: Uncontrolled Resource Consumption, CWE-89: SQL Injection, CWE-20: Improper Input Validation, CWE-22: Improper Limitation of a Pathname to a Restricted Directory, CWE-94: Code Injection, CWE-476: NULL Pointer Dereference, CWE-798: Use of Hard-coded Credentials, CWE-79: Cross-site Scripting, and CWE-416: Use After Free.
In the future, we plan to: (1) collect additional code snippets from other open source repositories and industrial projects, and code snippets generated by newer releases of Copilot; (2) analyze and summarize the application scenarios of these code snippets, studying how practitioners use Copilot and fix the issues in development; and (3) compare the results with other emerging Generative AI code generation tools such as CodeWhisperer, aiXcoder, and Code Llama.

### v2

@@ 5.2. Implications
Importance of continuous security analysis of Copilot-generated code. We conjecture practitioners using Copilot will likely encounter security weaknesses, regardless of the programming language used. Our study results reflect the inevitability of security weaknesses in Copilot-generated code, and there are a variety of security weaknesses in Copilot-generated code. Practitioners must be aware of the diverse security scenarios in production and adopt multiple security prevention measures to address security risks before accepting vulnerable code suggestions. Maintaining a rigorous process of security checks concurrently with code generation can identify potential vulnerability risks and rectify the security issues in time. Developers should follow the best practices for using code generation tools and always check the code suggestions generated by Copilot (or any code generation tools). For example, developers can establish a gated check-in build process for checking and preventing security issues when committing code generated by Copilot. Initially, we can turn to automated tools to continuously scan the Copilot-generated code for known security issues, such as CWEs. Recognizing that these automated tools may not always detect all security issues, especially newly emerging ones or those that are context-specific, it is necessary to conduct a subsequent manual assessment, including manual security code review for Copilot-generated code. By embracing this combined strategy for continuous security analysis, we can ensure a robust security shield for the code committing process with Copilot-generated code.
Prevention of security issues in Copilot-generated code. According to our study results, we provide the following suggestions on how to prevent potential security issues in Copilot-generated code: (1) Targeted security countermeasures: Based on the frequency of related CWEs in Copilot-generated code, practitioners can proactively prevent and address security issues in a targeted manner. When using Copilot to generate Python or JavaScript code, developers should focus on dedicated security weaknesses (CWEs) in Copilot-generated code of different programming languages (see Table 6). Furthermore, a recent study shows that security weaknesses appear in certain code suggestions, but not all (Majdinasab et al., 2024); consequently, developers should carefully select potentially more secure suggestions, with the assistance of tools, that do not expose the code to vulnerabilities. (2) Standardized security assessment: Common security weaknesses in software development are also prevalent in the code generated by Copilot. As a good practice, developers can use the CWE Top-25 list as a guide to understand which security weaknesses are most common and dangerous in the generated code and follow the mitigation measures for related CWEs provided by the MITRE (The MITRE Corporation, 2023b). Additionally, the CWE Top-25 provides a standardized approach for security assessment, and developers can also use it as a guide to perform security audits of the code generated by Copilot. (3) Enhancing prompt engineering to generate security code: The instruction tuning schemes in code generation not only impacts the utility of the code but also its security. By combining the security fine-tuning with standard instruction tuning, joint optimization of security and utility can be promoted (He et al., 2024). We should incorporate security considerations from the initial stages of code generation. For specific security scenarios that are prone to CWEs, we can improve the security of code by enhancing prompt engineering, e.g., with the prompt patterns proposed in (White et al., 2023).
@@ 6. Threats to Validity
Construct Validity: This study has three threats to construct validity: (1) Using keyword-based search – We used a keyword-based search to collect relevant code snippets from GitHub. The results obtained through the keyword-based search may not cover all the code snippets generated by Copilot on GitHub. We tried to mitigate this threat by constantly and iteratively refining the keywords and using synonyms. (2) Manual data filtering – We manually screened the results obtained from the keyword-based search by analyzing the comments, tags, and other metadata of the code snippets to determine whether they were generated by Copilot. Since this process was manually done, it may have been influenced by personal bias. In this regard, two authors conducted the pilot experiment independently to minimize the impact on the construct validity. (3) Manual association of CWEs – We manually associated the warning messages reported by the static analysis tools with a particular CWE, which may introduce personal subjective bias, threatening the construct validity. We employed two measures to mitigate this threat. First, since the list of CWEs is a tree structure with interconnections between them, we first matched the warning messages to a higher-level CWE, and further checked whether we can match the warning messages to a lower-level CWE with more specific definition. Second, to mitigate the personal bias, two authors independently assigned each security issue description a CWE ID. In case of disagreement, the two authors discussed it with the assessment by a third author (a security expert).
@@
External Validity: Our dataset consists of Copilot-generated code snippets collected from open-source projects on GitHub. During the filtering process, we excluded code that utilized Copilot to solve algorithmic problems, aiming to ensure that the collected data genuinely reflected real-world production environments. Since the data from GitHub are not diversified enough, we had a higher number of code snippets originating from the Game projects. This could result in a lack of comprehensiveness in the security scenarios involved. The peculiarity of the data source may make the dataset incomplete, thereby threatening the external validity of the results. Furthermore, we acknowledge the need to collect more diverse code snippets from different platforms to increase the generalizability of the results. We will consider adopting more diversified ways or platforms to collect code. Additionally, due to the limitations of static analysis tools themselves, these tools could not scan all CWEs, and there is a degree of false positives in the scan results (as the case with static analysis, in general, (Sui et al., 2020; Kang et al., 2022)). Although we used two widely used static analysis tools to increase the comprehensiveness of the scans and manually checked the results of the tool scans, the results may suffer from incompleteness.
@@
Reliability: We used multiple automated static analysis tools to analyze the Copilot-generated code snippets to improve security weaknesses detection. Developers have widely used these automated tools. The querying mechanism of these tools ensures that the scan results remain consistent when used multiple times. In addition, we performed two rounds of scanning with two tools for security checks on each code snippet, intending to complement the results of one tool with the other. By implementing these measures, we believe that our research results are reliable and these threats to reliability are mitigated.
@@ 7. Conclusions
Automatic code generation and recommendation have been an active research area due to the advancement of AI and, specifically, LLMs. AI code generation tools, such as Copilot, can significantly improve developers’ development efficiency, but can also introduce vulnerabilities and security risks. In this paper, we present the results of an empirical study to analyze security weaknesses in Copilot-generated code found in public GitHub projects. We identified 452 code snippets generated by Copilot from GitHub projects and analyzed those snippets for security weaknesses using static analysis tools. This study aims to help developers understand the security risks of weaknesses introduced in the code generated by Copilot (and potentially similar code generation tools). Our results show that (1) 29.8% of the 452 Copilot-generated code snippets contain security weaknesses. Developers have a high risk of raising security issues when using Copilot, regardless of the programming language, so security checks are necessary. (2) The detected security weaknesses are diverse in nature and are associated with 38 different CWEs. Developers face a variety of development scenarios and application environments in production and need the appropriate security awareness and skills. (3) Among these CWEs, eight appear in the MITRE CWE Top-25 list, and six belong to the Stubborn Weaknesses, demonstrating their severity.
In the future, we plan to: (1) collect additional code snippets from other open source repositories and industrial projects and code snippets generated by newer releases of Copilot; (2) analyze and summarize the application scenarios of these code snippets, studying how practitioners use Copilot and fix the issues in development; and (3) compare the results with other emerging Generative AI code generation tools such as CodeWhisperer, aiXcoder, and Code Llama.
Data availability
We have shared the link to our dataset in the reference (Fu et al., 2024).
Acknowledgments
This work is funded by the National Natural Science Foundation of China (NSFC) under Grant No. 62172311 and the Special Fund of Hubei Luojia Laboratory.